# Expected Information Gain Estimation via Density Approximations: Sample Allocation and Dimension Reduction

### v1

@@ Abstract
Computing expected information gain (EIG) from prior to posterior (equivalently, mutual information between candidate observations and model parameters or other quantities of interest) is a fundamental challenge in Bayesian optimal experimental design. We formulate flexible transport-based schemes for EIG estimation in general nonlinear/non-Gaussian settings, compatible with both standard and implicit Bayesian models. These schemes are representative of two-stage methods for estimating or bounding EIG using marginal and conditional density estimates. In this setting, we analyze the optimal allocation of samples between training (density estimation) and approximation of the outer prior expectation. We show that with this optimal sample allocation, the MSE of the resulting EIG estimator converges more quickly than that of a standard nested Monte Carlo scheme. We then address the estimation of EIG in high dimensions, by deriving gradient-based upper bounds on the mutual information lost by projecting the parameters and/or observations to lower-dimensional subspaces. Minimizing these upper bounds yields projectors and hence low-dimensional EIG approximations that outperform approximations obtained via other linear dimension reduction schemes. Numerical experiments on a PDE-constrained Bayesian inverse problem also illustrate a favorable trade-off between dimension truncation and the modeling of non-Gaussianity, when estimating EIG from finite samples in high dimensions.
@@ 1 Introduction
where $\operatorname{\mathcal{D}_{\text{KL}}}$ denotes the Kullback–Leibler (KL) divergence. This design objective was advocated by Lindley [41] as a natural measure of information due to an experiment. It can be justified from a decision-theoretic perspective [8, 15] as the expectation of a utility function based on Shannon information. The expected information gain enjoys other useful properties, such as invariance under reparameterization (inherited from the KL divergence). The EIG is also equivalent to the mutual information (MI) between $X$ and $Y|d$, denoted $\mathcal{I}(X;Y|d)$ [62]. We thus will use these two terms interchangeably. In the Bayesian linear-Gaussian model, maximizers of the expected information gain from prior to posterior coincide with Bayesian D-optimal designs, i.e., $d^{\ast}\in\operatorname*{arg\,min}_{d\in\Xi}\log\det\Sigma_{X|Y}(d)$, where $\Sigma_{X|Y}(d)$ is the design-dependent posterior covariance matrix. In general, however (e.g., outside of the linear-Gaussian setting), closed-form expressions for the EIG are unavailable. Estimating or bounding the EIG is in fact one of the central challenges of nonlinear Bayesian experimental design.
@@ 1.1 Related work
In the first setting, a workhorse approach to EIG estimation has been nested Monte Carlo (NMC), which to our knowledge was first proposed by Ryan [56]. The core idea is to independently estimate the evidence $\pi_{Y}(y)$ for different values of $y$ and to embed these estimates in an outer Monte Carlo estimator of the expectation over $Y$, following (4). NMC is biased at finite sample sizes, but asymptotically unbiased; it also converges more slowly than standard Monte Carlo (see Rainforth et al. [54] and Section 4). For more details on NMC and its variants, we refer to Huan et al. [31, Section 3.1]. In the third setting, a classical method for estimating MI from samples is the nonparametric Kraskov–Stögbauer–Grassberger (KSG) estimator [36], which is based on the statistics of nearest-neighbor distances. The error of this estimator scales poorly with dimension [26], however, and it does not take advantage of smoothness in the associated densities.
@@
A more modern family of sample-driven approaches involves constructing and optimizing variational bounds for EIG. In general, these approaches involve writing a variational representation of the mutual information and optimizing this representation (thus tightening the bound) over a family of so-called “critic” functions. Prominent examples include the ‘mutual information neural estimation’ approach of Belghazi et al. [7], which uses the Donsker–Varadhan variational representation of the KL divergence to construct a lower bound for MI, and then maximizes this bound over a family of critic functions represented by deep neural networks. Related methods include the “tractable unnormalized Barber–Agakov” bound introduced in Poole et al. [50] and the NWJ (Nguyen, Wainwright, Jordan) estimator of Nguyen et al. [45]. We refer to Poole et al. [50] for links among these and other variational bounds as well as trade-offs between the bias and variance of the resulting EIG estimators. One key point is that the optimal critic functions in these bounds typically coincide with the log-conditional density $\log\pi_{Y|X}$, the log-evidence $\log\pi_{Y}$ and/or the log-density ratio $\log\pi_{X|Y}-\log\pi_{X}=\log\pi_{Y|X}-\log\pi_{Y}$, depending on the specifics of each construction.
@@
The use of variational bounds is thus linked to density approximation, and an alternative approach is to optimize the deficit of a bound over families of density functions (i.e., non-negative functions that integrate to one) rather than generic critic functions. In the context of OED, this approach was first suggested by Foster et al. [25], who construct a lower bound for EIG by approximating the posterior density $\pi_{X|Y}$ and an upper bound for EIG by approximating the marginal $\pi_{Y}$. Foster et al. [25] demonstrate their approach using simple mean-field variational approximations in standard parametric families. Although such approximations are easy to implement, they do not in general yield tight bounds (e.g., if the true posterior or evidence is far from the chosen family). We also note that other estimators of EIG or mutual information are based on explicit density ratio estimation [53, 61, 9, 14]. In the machine learning literature, MI estimators that involve learning the density ratio are sometimes called ‘discriminative,’ while those based on learning the relevant densities directly are called ‘generative’ [40]. The comparative performance of these two classes of methods is in general not clear. Here, we will concentrate on the latter.
@@
To this end, recent work has replaced the simpler density approximations of Foster et al. [25] with more flexible density approximations based on transportation of measure. For instance, Orozco et al. [47], Dong et al. [22], Kennamer et al. [33] use conditional normalizing flows [55, 49] trained from samples of $\pi_{X,Y}$ to approximate the density of $\pi_{X|Y}$. Koval et al. [35] instead construct functional tensor-train approximations [18] of the triangular Knothe–Rosenblatt (KR) transport map using direct evaluations of the joint density $\pi_{X,Y}$. The approach we develop in this paper is also based on transport. In contrast with Koval et al. [35], we estimate the KR map from samples, and then use the associated plug-in estimator of any desired marginal and/or conditional density to form an EIG approximation. Depending on the problem setting, this approximation may also be an upper or lower bound. Our KR approximation is built on a general representation of monotone triangular maps that employs an invertible rectification operator, proposed in Baptista et al. [3]; this operator transforms generic (non-monotone) functions $f$ into monotone invertible functions, in particular triangular diffeomorphisms. The resulting EIG approximations are thus parameterized by $f$, and in this sense our EIG approximation comes full circle and is analogous to a critic-based variational bound. Crucially, under appropriate assumptions on the target distribution, we can take advantage of theory in Baptista et al. [3] that guarantees recovery of the exact KR rearrangement and hence the exact marginals and conditionals. Thus, as the class of functions chosen to represent $f$ is enriched, our bounds on the EIG can become arbitrarily tight. A specific version of EIG estimators explored in this paper was also demonstrated in Baptista et al. [4].
@@
We also note that most of the aforementioned EIG estimation methods, despite their widespread adoption, have only been applied in relatively low-dimensional regimes. Intuitively, these methods either explicitly or implicitly involve density estimation, and in the absence of further structure, the number of samples required for non-parametric or semi-parametric density estimation scales exponentially with dimension [43, 65]. Some papers on density ratio estimation do include high-dimensional examples, but these examples are either Gaussian [16] or the density ratio is parameterized by a simple exponential family [61]. Yet a recurring structure in high-dimensional Bayesian inference problems is that the change from prior to posterior is well-captured by a low-dimensional subspace of the parameter space [20, 72, 11, 19]. Similarly, conditioning on a low-dimensional projection of the realized observations $y$ might closely approximate the result of conditioning on the full vector $y$ [27, 32]. An information theoretic perspective on these projections is developed in Baptista et al. [2]. In the present paper, we will use this perspective to create dimension reduction methods for OED, with error guarantees on the estimated value of the EIG.
Prior work on dimension reduction for OED has taken several forms. [69] uses Laplace approximations of the posterior, averaged over many realizations of the data, to approximate the EIG. These Laplace approximations are simplified using low-rank approximations of the prior-preconditioned Hessian of the log-likelihood, which is equivalent to parameter-space dimension reduction for linear (or linearized) problems [59]. Alternatively, in the setting of nonlinear inverse problems with additive noise, e.g., $Y=G(x)+\mathcal{E}$Wu et al. [70] constructs deep neural network approximations of the deterministic parameter-to-observable map $G$ after performing dimension reduction in both the input space (of parameters $X$) and output space (of observations $Y$). The resulting surrogate is then used in an NMC estimator of the EIG. We employ the same parameter-space projection here, but a different projection of the observations that yields sharper error bounds. Moreover, instead of using NMC, we use both dimension reduction methods in transport-based EIG estimators.
@@ 1.2 Contributions
As our first technical contribution, we then analyze the underlying problem of sample allocation and convergence rates for such EIG estimators. Our results apply to any parametric density estimation scheme used within the EIG estimation framework. We develop an asymptotically optimal allocation of samples, between density estimation and the estimation of the outer expectation, to balance the decay rates of bias and variance and thus maximize the convergence rate of the mean squared error (MSE) of the EIG estimate. With this allocation, we show that this class of estimators converges at a faster rate than an optimally-balanced NMC scheme. We numerically verify the conditions yielding optimal rates, in a simple linear-Gaussian setting where the exact EIG is available analytically. As a demonstration, we then evaluate the performance and versatility of density-based EIG estimation algorithms for a non-Gaussian problem with various EIG objectives.
Our second technical contribution is to combine dimension reduction with transport-based EIG estimation to construct estimators of EIG in high dimensions. Our proposed dimension reduction scheme relies on minimizing an upper bound for the error in EIG induced by projecting both the parameters $X$ and observations $Y$ to lower-dimensional subspaces. The scheme for computing these projections originates in [2], where it was used for the purpose of posterior approximation, but this information theoretic perspective on dimension reduction is a natural fit for EIG in non-Gaussian settings. It seeks $(r\ll n_{x})$-dimensional and $(s\ll n_{y})$-dimensional projections $X_{r}$ and $Y_{s}$ of $X$ and $Y$, respectively, that minimize an upper bound for the information loss $\mathcal{I}(X;Y)-\mathcal{I}(X_{r};Y_{s})$ at any $r$ and $s$. The preceding EIG estimation schemes can then be applied to the joint distribution of $X_{r}$ and $Y_{s}$. We show that this linear dimension reduction approach outperforms other linear approaches, such as principal component analysis (PCA) and canonical correlation analysis (CCA). We also demonstrate how this scheme allows one to trade off dimensionality with the modeling of non-Gaussianity: in many EIG estimation problems it is better to truncate the representation of the parameters and observations in a suitable basis and capture the non-Gaussianity of the resulting $\pi_{X_{r},Y_{s}}$, than it is to make cruder (e.g., Gaussian) approximations of the full distribution $\pi_{X,Y}$.
@@ 2.1 EIG bounds and estimators
Note that if we instead substitute approximations for the prior or likelihood in (3) or (4), EIG can neither be upper- nor lower-bounded [25]. That is, neither $\mathbb{E}_{\pi_{X,Y}}[\log\pi_{X|Y}(X|Y)-\log\widetilde{\pi}_{X}(X)]$ nor $\mathbb{E}_{\pi_{X,Y}}[\log\widetilde{\pi}_{Y|X}(Y|X)-\log\pi_{Y}(Y)]$ is a bound for EIG. Moreover, in certain problem settings—for instance, with an intractable likelihood and prior—one must approximate both the density in the numerator and the density in the denominator, and thus use Monte Carlo estimators of $\mathbb{E}_{\pi_{X,Y}}\left[\log\frac{\widetilde{\pi}_{X|Y}(x|y)}{\widetilde{\pi}_{X}(x)}\right]$ and $\mathbb{E}_{\pi_{X,Y}}\left[\log\frac{\widetilde{\pi}_{Y|X}(y|x)}{\widetilde{\pi}_{Y}(y)}\right]$. While these estimators can be consistent, they do not provide (stochastic) bounds for EIG.
@@ 2.2 Transport maps for density estimation
Measure transport provides an expressive and infinitely refinable way of approximating complex probability densities. Here we give a brief and intuitive introduction to this approach; for more details, we refer the reader to Marzouk et al. [42], Spantini et al. [60], Baptista et al. [3]. To illustrate key ideas, we first set aside the OED problem and the specific probability distributions it entails, returning to them in Section 2.2.3.
@@ 2.2.1 Maximum likelihood estimation of transport maps
where $\{z^{i}\}_{i=1}^{N}$ are i.i.d. samples from $\pi$. Let $\widehat{S}^{k}_{\mathrm{OPT}}=\operatorname*{arg\,min}_{S^{k}\in\mathcal{S}^{k}}\widehat{J}_{k}(S^{k})$, where $\mathcal{S}^{k}$ is a space of maps from $\mathbb{R}^{k}$ to $\mathbb{R}$ satisfying the monotonicity constraint $\partial_{k}S^{k}>0$. As explained in Baptista et al. [3], Wang and Marzouk [65], each $\widehat{S}^{k}_{\mathrm{OPT}}$ is a maximum likelihood estimate of the corresponding map component $S^{k}$, and concatenating these functions yields a maximum likelihood estimate of the entire triangular map: $\widehat{S}_{\mathrm{OPT}}=(\widehat{S}^{1}_{\mathrm{OPT}},\widehat{S}^{2}_{\mathrm{OPT}},\ldots,\widehat{S}^{n}_{\mathrm{OPT}})$.

### v2

@@ Abstract
Computing expected information gain (EIG) from prior to posterior (equivalently, mutual information between candidate observations and model parameters or other quantities of interest) is a fundamental challenge in Bayesian optimal experimental design. We formulate flexible transport-based schemes for EIG estimation in general nonlinear/non-Gaussian settings, compatible with both standard and implicit Bayesian models. These schemes are representative of two-stage methods for estimating or bounding EIG using marginal and conditional density estimates. In this setting, we analyze the optimal allocation of samples between training (density estimation) and approximation of the outer prior expectation. We show that with this optimal sample allocation, the mean squared error (MSE) of the resulting EIG estimator converges more quickly than that of a standard nested Monte Carlo scheme. We then address the estimation of EIG in high dimensions, by deriving gradient-based upper bounds on the mutual information lost by projecting the parameters and/or observations to lower-dimensional subspaces. Minimizing these upper bounds yields projectors and hence low-dimensional EIG approximations that outperform approximations obtained via other linear dimension reduction schemes. Numerical experiments on a PDE-constrained Bayesian inverse problem also illustrate a favorable trade-off between dimension truncation and the modeling of non-Gaussianity, when estimating EIG from finite samples in high dimensions.
@@ 1 Introduction
where $\operatorname{\mathcal{D}_{\text{KL}}}$ denotes the Kullback–Leibler (KL) divergence. This design objective was advocated by [41] as a natural measure of information due to an experiment. It can be justified from a decision-theoretic perspective [8, 15] as the expectation of a utility function based on Shannon information. The expected information gain enjoys other useful properties, such as invariance under reparameterization (inherited from the KL divergence). The EIG is also equivalent to the mutual information (MI) between $X$ and $Y|d$, denoted $\mathcal{I}(X;Y|d)$ [62]. We thus will use these two terms interchangeably. In the Bayesian linear-Gaussian model, maximizers of the expected information gain from prior to posterior coincide with Bayesian D-optimal designs, i.e., $d^{\ast}\in\operatorname*{arg\,min}_{d\in\Xi}\log\det\Sigma_{X|Y}(d)$, where $\Sigma_{X|Y}(d)$ is the design-dependent posterior covariance matrix. In general, however (e.g., outside of the linear-Gaussian setting), closed-form expressions for the EIG are unavailable. Estimating or bounding the EIG is in fact one of the central challenges of nonlinear Bayesian experimental design.
@@ 1.1 Related work
In the first setting, a workhorse approach to EIG estimation has been nested Monte Carlo (NMC), which to our knowledge was first proposed by [56]. The core idea is to independently estimate the evidence $\pi_{Y}(y)$ for different values of $y$ and to embed these estimates in an outer Monte Carlo estimator of the expectation over $Y$, following (4). NMC is biased at finite sample sizes, but asymptotically unbiased; it also converges more slowly than standard Monte Carlo (see [54] and Section 4). For more details on NMC and its variants, we refer to [32, Section 3.1]. In the third setting, a classical method for estimating MI from samples is the nonparametric Kraskov–Stögbauer–Grassberger (KSG) estimator [37], which is based on the statistics of nearest-neighbor distances. The error of this estimator scales poorly with dimension [27], however, and it does not take advantage of smoothness in the associated densities.
@@
A more modern family of sample-driven approaches involves constructing and optimizing variational bounds for EIG. In general, these approaches involve writing a variational representation of the mutual information and optimizing this representation (thus tightening the bound) over a family of so-called “critic” functions. Prominent examples include the ‘mutual information neural estimation’ approach of [7], which uses the Donsker–Varadhan variational representation of the KL divergence to construct a lower bound for MI, and then maximizes this bound over a family of critic functions represented by deep neural networks. Related methods include the “tractable unnormalized Barber–Agakov” bound introduced in [50] and the NWJ (Nguyen, Wainwright, Jordan) estimator of [45]. We refer to [50] for links among these and other variational bounds as well as trade-offs between the bias and variance of the resulting EIG estimators. One key point is that the optimal critic functions in these bounds typically coincide with the log-conditional density $\log\pi_{Y|X}$, the log-evidence $\log\pi_{Y}$ and/or the log-density ratio $\log\pi_{X|Y}-\log\pi_{X}=\log\pi_{Y|X}-\log\pi_{Y}$, depending on the specifics of each construction.
@@
The use of variational bounds is thus linked to density approximation, and an alternative approach is to optimize the deficit of a bound over families of density functions (i.e., non-negative functions that integrate to one) rather than generic critic functions. In the context of OED, this approach was first suggested by [26], who construct a lower bound for EIG by approximating the posterior density $\pi_{X|Y}$ and an upper bound for EIG by approximating the marginal $\pi_{Y}$. [26] demonstrate their approach using simple mean-field variational approximations in standard parametric families. Although such approximations are easy to implement, they do not, in general, yield tight bounds (e.g., if the true posterior or evidence is far from the chosen family). We also note that other estimators of EIG or mutual information are based on explicit density ratio estimation [53, 61, 9, 14]. In the machine learning literature, MI estimators that involve learning the density ratio are sometimes called ‘discriminative,’ while those based on learning the relevant densities directly are called ‘generative’ [40]. The comparative performance of these two classes of methods is in general not clear. Here, we will concentrate on the latter.
@@
To this end, recent work has replaced the simpler density approximations of [26] with more flexible density approximations based on transportation of measure. For instance, [47, 23, 34] use conditional normalizing flows [55, 49] trained from samples of $\pi_{X,Y}$ to approximate the density of $\pi_{X|Y}$. [36] instead constructs functional tensor-train approximations [18] of the triangular Knothe–Rosenblatt (KR) transport maps using direct evaluations of the joint density $\pi_{X,Y}$, which is further extended by [19] (combined with dimension reduction) to the sequential design setting. The approach we develop in this paper is also based on transport. In contrast with [36], we estimate the KR map from samples, and then use the associated plug-in estimator of any desired marginal and/or conditional density to form an EIG approximation. Depending on the problem setting, this approximation may also be an upper or lower bound. Our KR approximation is built on a general representation of monotone triangular maps that employs an invertible rectification operator, proposed in [5]; this operator transforms generic (non-monotone) functions $f$ into monotone invertible functions, in particular triangular diffeomorphisms. The resulting EIG approximations are thus parameterized by $f$, and in this sense our EIG approximation comes full circle and is analogous to a critic-based variational bound. Crucially, under appropriate assumptions on the target distribution, we can take advantage of theory in [5] that guarantees recovery of the exact KR rearrangement and hence the exact marginals and conditionals. Thus, as the class of functions chosen to represent $f$ is enriched, our bounds on the EIG can become arbitrarily tight. A specific version of EIG estimators explored in this paper was also demonstrated in [2].
@@
We also note that most of the aforementioned EIG estimation methods, despite their widespread adoption, have only been applied in relatively low-dimensional regimes. Intuitively, these methods either explicitly or implicitly involve density estimation, and in the absence of further structure, the number of samples required for non-parametric or semi-parametric density estimation scales exponentially with dimension [43, 66]. Some papers on density ratio estimation do include high-dimensional examples, but these examples are either Gaussian [16] or the density ratio is parameterized by a simple exponential family [61]. Yet a recurring structure in high-dimensional Bayesian inference problems is that the change from prior to posterior is well-captured by a low-dimensional subspace of the parameter space [20, 73, 11, 21, 19]. Similarly, conditioning on a low-dimensional projection of the realized observations $y$ might closely approximate the result of conditioning on the full vector $y$ [28, 33]. An information theoretic perspective on these projections is developed in [4]. In the present paper, we will use this perspective to create dimension reduction methods for OED, with error guarantees on the estimated value of the EIG.
Prior work on dimension reduction for OED has taken several forms. [70] uses Laplace approximations of the posterior, averaged over many realizations of the data, to approximate the EIG. These Laplace approximations are simplified using low-rank approximations of the prior-preconditioned Hessian of the log-likelihood, which is equivalent to parameter-space dimension reduction for linear (or linearized) problems [60]. [19] utilizes a likelihood-informed subspace to identify the informative direction, which is subsequently used to reduce the dimensionality of the parameters. Alternatively, in the setting of nonlinear inverse problems with additive noise, e.g., $Y=G(x)+\mathcal{E}$[71] constructs deep neural network approximations of the deterministic parameter-to-observable map $G$ after performing dimension reduction in both the input space (of parameters $X$) and output space (of observations $Y$). The resulting surrogate is then used in an NMC estimator of the EIG. We employ the same parameter-space projection here, but a different projection of the observations that yields sharper error bounds. Moreover, instead of using NMC, we use both dimension reduction methods in transport-based EIG estimators.
@@ 1.2 Contributions
As our first technical contribution, we analyze the underlying problem of sample allocation and convergence rates for such EIG estimators. Our results apply to any parametric density estimation scheme used within the EIG estimation framework. We develop an asymptotically optimal allocation of samples, between density estimation and the estimation of the outer expectation, to balance the decay rates of bias and variance and thus maximize the convergence rate of the mean squared error (MSE) of the EIG estimate. With this allocation, we show that this class of estimators converges at a faster rate than an optimally-balanced NMC scheme. We numerically verify the conditions yielding optimal rates, in a simple linear-Gaussian setting where the exact EIG is available analytically. As a demonstration, we then evaluate the performance and versatility of density-based EIG estimation algorithms for a non-Gaussian problem with various EIG objectives.
Our second technical contribution is to combine dimension reduction with transport-based EIG estimation to construct estimators of EIG in high dimensions. Our proposed dimension reduction scheme relies on minimizing an upper bound for the error in EIG induced by projecting both the parameters $X$ and observations $Y$ to lower-dimensional subspaces. The scheme for computing these projections originates in [4], where it was used for the purpose of posterior approximation, but this information theoretic perspective on dimension reduction is a natural fit for EIG in non-Gaussian settings. It seeks $(r\ll n_{x})$-dimensional and $(s\ll n_{y})$-dimensional projections $X_{r}$ and $Y_{s}$ of $X$ and $Y$, respectively, that minimize an upper bound for the information loss $\mathcal{I}(X;Y)-\mathcal{I}(X_{r};Y_{s})$ at any $r$ and $s$. The preceding EIG estimation schemes can then be applied to the joint distribution of $X_{r}$ and $Y_{s}$. We show that this dimension reduction approach outperforms other linear approaches, such as principal component analysis (PCA) and canonical correlation analysis (CCA). We also demonstrate how this scheme allows one to trade off dimensionality with the modeling of non-Gaussianity: in many EIG estimation problems it is better to truncate the representation of the parameters and observations in a suitable basis and capture the non-Gaussianity of the resulting $\pi_{X_{r},Y_{s}}$, than it is to make cruder (e.g., Gaussian) approximations of the full distribution $\pi_{X,Y}$.
@@ 2.1 EIG bounds and estimators
If we replace the prior or likelihood in equations (3) or (4) with approximations $\widetilde{\pi}_{X}$ and $\widetilde{\pi}_{Y|X}$, the expected information gain (EIG) is neither upper-bounded nor lower-bounded [26]. Specifically, neither $\mathbb{E}_{\pi_{X,Y}}\left[\log\frac{\widetilde{\pi}_{X|Y}(x|y)}{\widetilde{\pi}_{X}(x)}\right]$ nor $\mathbb{E}_{\pi_{X,Y}}\left[\log\frac{\widetilde{\pi}_{Y|X}(y|x)}{\widetilde{\pi}_{Y}(y)}\right]$ serves as a bound for EIG. Furthermore, in cases where both the prior and likelihood are intractable, it becomes necessary to approximate both the numerator and denominator densities. This requires using Monte Carlo estimators for $\mathbb{E}_{\pi_{X,Y}}\left[\log\frac{\widetilde{\pi}_{X|Y}(x|y)}{\widetilde{\pi}_{X}(x)}\right]$ and $\mathbb{E}_{\pi_{X,Y}}\left[\log\frac{\widetilde{\pi}_{Y|X}(y|x)}{\widetilde{\pi}_{Y}(y)}\right]$. While these estimators can be consistent, they do not provide (stochastic) bounds for EIG.
@@ 2.2 Transport maps for density estimation
Measure transport provides an expressive and infinitely refinable way of approximating complex probability densities. Here we give a brief and intuitive introduction to this approach; for more details, we refer the reader to [42, 59, 5]. To illustrate key ideas, we first set aside the OED problem and the specific probability distributions it entails, returning to them in Section 2.2.3.
@@ 2.2.1 Maximum likelihood estimation of transport maps
where $\{z^{i}\}_{i=1}^{N}$ are i.i.d. samples from $\pi$. Let $\widehat{S}^{k}_{\mathrm{OPT}}=\operatorname*{arg\,min}_{S^{k}\in\mathcal{S}^{k}}\widehat{J}_{k}(S^{k})$, where $\mathcal{S}^{k}$ is a space of maps from $\mathbb{R}^{k}$ to $\mathbb{R}$ satisfying the monotonicity constraint $\partial_{k}S^{k}>0$. As explained in [5, 66], each $\widehat{S}^{k}_{\mathrm{OPT}}$ is a maximum likelihood estimate of the corresponding map component $S^{k}$, and concatenating these functions yields a maximum likelihood estimate of the entire triangular map: $\widehat{S}_{\mathrm{OPT}}=(\widehat{S}^{1}_{\mathrm{OPT}},\widehat{S}^{2}_{\mathrm{OPT}},\ldots,\widehat{S}^{n}_{\mathrm{OPT}})$.